# ConLID: Supervised Contrastive Learning for Low-Resource Language Identification

### v1

@@ Abstract
Language identification (LID) is a critical step in curating multilingual LLM pretraining corpora from web crawls. While many studies on LID model training focus on collecting diverse training data to improve performance, low-resource languages – often limited to single-domain data, such as the Bible – continue to perform poorly. To resolve these class imbalance and bias issues, we propose a novel supervised contrastive learning (SCL) approach to learn domain-invariant representations for low-resource languages. Through an extensive analysis, we show that our approach improves LID performance on out-of-domain data for low-resource languages by 3.2%, demonstrating its effectiveness in enhancing LID models. [^1] Our code and model are available at https://github.com/epfl-nlp/ConLID.
@@
Negar Foroutan1thanks: Equal contribution, Jakhongir Saydaliev [^1] footnotemark: 1, Ye Eun Kim2, Antoine Bosselut1 1EPFL  2The University of Texas at Austin Correspondence: {negar.foroutan, antoine.bosselut}@epfl.ch
@@ 1 Introduction
Figure 1: Overview of training ConLID. Each sentence is first processed by an encoder that generates a sentence representation based on the character-level n-grams of its constituent words. This representation is then passed to two components: (a) a Supervised Contrastive Learning (SCL) module, where positive and negative pairs are sampled using either soft or hard sampling strategies, and (b) a feed-forward neural network that serves as the classification head. The final loss used for training is a combination of the classification loss and the SCL loss.
@@
The application of these LID systems to web crawls requires scalable methods that can be efficiently applied to large document collections (Penedo et al., 2024). Consequently, current approaches to LID (Kargaran et al., 2023; Team et al., 2022) predominantly use simple models (Joulin et al., 2017) trained with cross-entropy (CE) loss. While these supervised approaches have been sufficient to practically solve LID for high-resource standard languages (McNamee, 2005), low-resource languages and dialects remain difficult to identify and categorize accurately for two reasons. First, data in these languages is often scarce or even mislabeled (Roy et al., 2022; Yu et al., 2022; Joshi et al., 2020), leading to class imbalances during training (Kargaran et al., 2023). Second, available data is often concentrated in specific domains, such as religious texts (i.e., Bible translations; Agić and Vulić, 2019; Imani et al., 2023; Ebrahimi and Kann, 2021). This domain entanglement yields models trained on narrow datasets that often fail to generalize effectively to diverse types of text.
@@
To mitigate these limitations, we propose a novel approach that employs supervised contrastive learning (SCL; Khosla et al., 2020) for LID. Our SCL objective explicitly encourages representations of texts from the same language to cluster together while pushing those from different languages further apart in the embedding space. Combined with a classic CE loss, our dual-objective framework promotes learning less biased language representations that are robust to in-language domain shifts, enabling better generalization across diverse text types. A critical component of our SCL training method is a memory bank Tan et al. (2022) from which we scalably sample more diverse positive and negative examples, effectively increasing our training batch size while keeping memory requirements constant. This augmentation enhances our contrastive learning approach, yielding stable training and discriminative language representations. We also employ a novel hard negative mining scheme for SCL training, where data sources in our training set are labeled with domain information and negative samples are drawn from different languages in the same domain to learn language-specific, domain-invariant representations during training.
@@
We evaluate our approach (called ConLID) on three benchmark datasets: GlotLID-C (Kargaran et al., 2023), FLORES-200 Team et al. (2022), and UDHR, [^2] https://huggingface.co/datasets/cis-lmu/udhr-lid and demonstrate that our approach leads to improvements over state-of-the-art LID methods, particularly in out-of-domain evaluations. Specifically, our best SCL-based approaches achieve a 3.2% improvement over CE-based models for low-resource languages, and a 5.4% improvement for languages with data from diverse domains, indicating enhanced domain generalization.
@@
Moreover, we conduct an in-depth analysis of underperforming languages in out-of-domain evaluations and observe that misclassifications predominantly occur among linguistically related languages, highlighting potential challenges in LID for closely related language pairs. Finally, we assess ConLID in a real-world scenario by evaluating it on FineWeb-2 (Penedo et al., 2024), a large-scale multilingual pretraining corpus. We compare its performance against state-of-the-art LID systems for low-resource languages, providing insights into its effectiveness in practical applications.
@@ 3 Methodology
In the following, we provide a detailed description of our base model architecture (§3.1), the principles of supervised contrastive learning (SCL; §3.2), our positive/negative pair selection approach (§3.3), and introduce an optimized training strategy for the SCL objective, particularly suited for scenarios involving a large number of classes (§3.4).
@@ 3.1 Base Classifier
We train the base model following the methodology employed by FastText (Joulin et al., 2017). Given an input sentence, each word in the sentence is decomposed into character-level n-grams. The sentence representation is obtained by computing the average of the vectors corresponding to these n-grams, along with the word-level embedding. The training process proceeds as follows:
Initially, the training dataset is partitioned into mini-batches. For a sampled mini-batch of size $B$, we define $S=\left\{x_{i},y_{i}\right\}_{i=1}^{B}$, where $x_{i}$ represents the input text, and $y_{i}$ denotes its corresponding language class. Each input text $x_{i}$ is represented by an embedding $\mathbf{z}_{i}$ using an encoder function $E$; $\mathbf{z}_{i}=E(x_{i})$. The resulting embeddings $\mathbf{z}_{i}$ are subsequently passed through a feed-forward neural network $f$, which serves as the classification head. The model is trained using the cross-entropy (CE) loss, formulated as:
@@ 3.3 Negative Pairs Selection
Given a batch $S=\left\{x_{i},y_{i};s_{i},d_{i}\right\}_{i=1}^{B}$, where $s_{i}$ and $d_{i}$ denote the script and domain of $x_{i}$, negative pairs are chosen according to Algorithm 1. For each sample $x_{i}$ in the batch, we select negative pair samples with a different label $y$, the same script $s$, and the same domain $d$. If fewer than $K$ such negatives exist, we relax the conditions in the following order: (2) different label $y$ and same script $s$, (3) different label $y$ and same domain $d$, (4) different label $y$ only (i.e., equivalent to Soft Selection).
Algorithm 1 Hard Negative Sample Selection
1:  for $i<B$ do
3:     if $|N(i)|<K$ then
4:        $N(i)=\left\{z_{j}\in S,y_{j}\neq y_{i},s_{j}=s_{i}\right\}$Condition 2}
5:     end if
6:     if $|N(i)|<K$ then
7:        $N(i)=\left\{z_{j}\in S,y_{j}\neq y_{i},d_{j}=d_{i}\right\}$Condition 3}
8:     end if
9:     if $|N(i)|<K$ then
10:        $N(i)=\left\{z_{j}\in S,y_{j}\neq y_{i}\right\}$Condition 4}
11:     end if
12:  end for
13:  return  $N(i)$negative pairs for each sample}
@@ 4.1 Training
(a) $\text{LID}_{\text{CE}}$ The model trained only with the cross-entropy objective, as defined in Equation (1).
(b) $\text{LID}_{\text{SCL}}$ The model trained with the combined objective in Equation (3), without employing a memory bank.
@@
In our experiments, we implemented $\text{LID}_{\text{CE}}$ as a replication of GlotLID-M Kargaran et al. (2023), as official data splits for GlotLID-M were not openly released (hindering a non-contaminated evaluation on GlotLID-C-test). However, to enable a full comparison with prior methods, we also report the performance of the open-weight GlotLID-M, the previous SOTA model on UDHR, and GlotLID-M+ConLID-S, our ensemble of GlotLID-M and ConLID-S by taking the average of their predicted distributions. Technical details can be found in Appendix B.1.
@@ 4.2 Evaluation
$\text{LID}_{\text{CE}}$98586e-4 .00000681.2e-7 .892519e-4 .000222234e-7 .96503e-4 .00011429.9e-7
$\text{LID}_{\text{SCL}}$98331e-4 .00000750.2e-7 .893819e-4 .000209518e-7 .96173e-4 .000118225e-7
@@
The original UDHR dataset contains over 500 translations of the Universal Declaration of Human Rights. Kargaran et al. (2023) curated and refined this dataset, reducing it to 369 languages. During our experiments, we identified discrepancies in seven languages (dip_Latn, pcd_Latn, aii_Syrc, guu_Latn, mto_Latn, chj_Latn, auc_Latn) due to character misalignments between GlotLID-C and UDHR, and two languages (taj_Deva, tzm_Tfng) that had only one example each. To ensure consistency, we exclude these nine languages, resulting in a final UDHR dataset comprising 360 languages. UDHR serves as an out-of-domain evaluation dataset and is not included in GlotLID-C-train-down.
@@
This dataset is a multilingual pretraining corpus encompassing $\sim$1,800 languages (Penedo et al., 2024). The LID process in the FineWeb-2 pipeline is conducted using GlotLID-M Kargaran et al. (2023), a model capable of recognizing 2102 languages. We use this dataset to evaluate our model in large-scale real-world scenarios, which provides insights into the effectiveness of our approach for LID for web crawls.
@@
Resource Level $\text{LID}_{\text{CE}}$ $\Delta$$\text{LID}_{\text{SCL}}$ $\Delta$ConLID-S $\Delta$ConLID-S+LIDCE
High-resource 0.9065 0 0.0066 0.0066
Low-resource 0.7380 0.0168 0.0323 0.0154
Table 2: F1 score change on UDHR of the LID models with SCL stratified by language resource level. $\text{LID}_{\text{CE}}$ is used as a baseline, and subtracted from each model’s score to compute the $\Delta$ score, e.g., $\Delta$$\text{LID}_{\text{SCL}}$= $\text{LID}_{\text{SCL}}$-$\text{LID}_{\text{CE}}$.
Domains $\text{LID}_{\text{CE}}$ $\Delta$$\text{LID}_{\text{SCL}}$ $\Delta$ConLID-S $\Delta$ConLID-S+LIDCE
Random 0.8157 0.0109 0.0541 0.0329
Bible 0.7821 -0.0077 0.0190 0.0149
Table 3: F1 score change of the LID models’ with SCL by GlotLID-C train domains. $\text{LID}_{\text{CE}}$ is used as a baseline, and subtracted from each model’s score to compute the absolute $\Delta$ score, e.g., $\Delta$$\text{LID}_{\text{SCL}}$= $\text{LID}_{\text{SCL}}$-$\text{LID}_{\text{CE}}$.
@@ 5 Experimental Analysis
Table 1 presents the overall results for GlotLID-C-test, UDHR, and FLORES-200. We observe no significant improvement when directly applying the SCL objective ($\text{LID}_{\text{SCL}}$) compared to the cross-entropy baseline ($\text{LID}_{\text{CE}}$). However, when incorporating a memory bank to increase the number of contrastive negatives beyond the number of classes, we achieve performance gains across all three evaluation datasets. Notably, this improvement is most significant for UDHR (out-of-domain), underscoring the effectiveness of contrastive learning in enhancing generalization.
For in-domain evaluation (GlotLID-C-test and FLORES-200), the highest F1 score among our methods is an ensemble approach that combines $\text{LID}_{\text{CE}}$ and ConLID-S with a memory bank of $\text{M}_{\text{H}}$=2048 and K=1024, suggesting that LIDCE already performs well for in-domain scenarios, while ConLID-S+LIDCE effectively balances in-domain performance with improved generalization.
When comparing to prior at, we observe that GlotLID-M tends to outperform our methods on FLORES-200, but underperforms them on UDHR, which is slightly more out of distribution. The ensembled method (GlotLID-M+ConLID-S), however, achieves the best score across UDHR and FLORES-200, indicating the methods are complementary and could be combined for optimal accuracy (at the cost of computational cost). No comparison is made on GlotLID-C-test as GlotLID-M was trained on different subsets of GlotLID-C, meaning the evaluation would be contaminated (same for GlotLID-M+ConLID-S).
@@ 5.1 Performance on UDHR Dataset
Script $\text{LID}_{\text{CE}}$ $\Delta$$\text{LID}_{\text{SCL}}$ $\Delta$ConLID-S $\Delta$ConLID-S+LIDCE
@@
Table 4: F1 score change of the LID models’ with SCL by language script. $\text{LID}_{\text{CE}}$ is used as a baseline, and subtracted from each model’s score to compute the absolute $\Delta$ score, e.g., $\Delta$$\text{LID}_{\text{SCL}}$= $\text{LID}_{\text{SCL}}$-$\text{LID}_{\text{CE}}$.
Resource Analysis.
We categorize languages into high- and low-resource groups based on the number of training examples available in GlotLID-C-train-down. Languages with fewer than 10k examples are designated as low-resource, while those with 10k or more examples are categorized as high-resource.
For each group of languages, we compute the average F1 score difference between SCL-based models ($\text{LID}_{\text{SCL}}$, ConLID-S, ConLID-S+LIDCE) and $\text{LID}_{\text{CE}}$. The results are presented in Table 2. Our analysis indicates that the model ConLID-S yields the most substantial improvements for low-resource languages, with F1 score gain of 3.23%, validating the importance of memory bank (§3.4). In contrast, the performance gain for high-resource languages is more modest, at 0.66%.
Domain Analysis.
The GlotLID-C dataset comprises five distinct textual domains: Bible Translations, Literature (i.e., stories & books), Politics (i.e., government documents), Multi-domain (News and Wikipedia articles), and Random (machine translation datasets, chat logs, and web-scraped data). To evaluate the impact of SCL across different training data domains, we analyze languages trained exclusively on a single domain. While the distribution of GlotLID-C-train-down is shown in Appendix A, among the languages in UDHR, 7 languages have been exclusively trained on Random, while 95 on Bible. Therefore, we compute the F1 score difference between SCL-based models ($\text{LID}_{\text{SCL}}$, ConLID-S, ConLID-S+LIDCE) and $\text{LID}_{\text{CE}}$ for random and bible domains and present the results in Table 3. Our results indicate that the inclusion of a memory bank leads to performance improvements across both domains. Notably, the Random datasets exhibit a more significant gain, with improvements of up to 5.41%. We hypothesize that this is due to the broader range of subdomains covered by Random, which facilitate learning more generalized language representations. In contrast, the Bible domain is limited to religious texts, making it less conducive to generalization.
Script Analysis.
We conduct a similar analysis on the scripts of the languages and present the results in Table 4. While the UDHR dataset comprises 31 scripts, we report only those for which at least one of the metrics $\Delta$$\text{LID}_{\text{SCL}}$, $\Delta$ConLID-S, $\Delta$ConLID-S+LIDCE is nonzero. As observed, ConLID-S exhibits an improvement of up to 10.9% (on Java), while some high-resource scripts show non-significant improvement (e.g., Hani, Cyrillic). This can be attributed to the resource level of these scripts, which aligns with our findings from §5.1, particularly regarding the behavior of low-resource languages. A breakdown analysis can be found in the Appendix C.2.

### v2

@@ Abstract
Language identification (LID) is a critical step in curating multilingual LLM pretraining corpora from web crawls. While many studies on LID model training focus on collecting diverse training data to improve performance, low-resource languages – often limited to single-domain data, such as the Bible – continue to perform poorly. To resolve these imbalance and bias issues, we propose a novel supervised contrastive learning (SCL) approach to learn domain-invariant representations for low-resource languages. We show that our approach improves LID performance on out-of-domain data for low-resource languages by 3.2 percentage points, while maintaining its performance for the high-resource languages. [^1] Our code and model are available at https://github.com/epfl-nlp/ConLID.
@@
Negar Foroutan1thanks: Equal contribution, Jakhongir Saydaliev [^1] footnotemark: 1, Grace Kim2, Antoine Bosselut1 1EPFL  2The University of Texas at Austin Correspondence: {negar.foroutan, antoine.bosselut}@epfl.ch
@@ 1 Introduction
Figure 1: Overview of training ConLID. Each sentence is first processed by an encoder that generates a sentence representation. This representation is then passed to two components: (a) a Supervised Contrastive Learning (SCL) module, and (b) a feed-forward neural network that serves as the classification head. The final loss used for training is a combination of the classification loss and the SCL loss.
@@
The application of these LID systems to web crawls requires scalable methods that can be efficiently applied to large document collections (Penedo et al., 2025). Consequently, current approaches to LID (Kargaran et al., 2023; Team et al., 2022) predominantly use simple models (Joulin et al., 2017) trained with cross-entropy (CE) loss. While these supervised approaches have been sufficient to practically solve LID for high-resource standard languages (McNamee, 2005), low-resource languages and dialects remain difficult to identify and categorize accurately for two reasons. First, data in these languages is often scarce or even mislabeled (Roy et al., 2022; Yu et al., 2022; Joshi et al., 2020a), leading to class imbalances during training (Kargaran et al., 2023). Second, available data is often concentrated in specific domains, such as religious texts (i.e., Bible translations; Agić and Vulić, 2019; Imani et al., 2023; Ebrahimi and Kann, 2021). This domain entanglement yields models trained on narrow datasets that often fail to generalize effectively to diverse types of text.
@@
To mitigate these limitations, we propose a novel approach that employs supervised contrastive learning (SCL; Khosla et al., 2020) for LID. Our SCL objective explicitly encourages representations of texts from the same language to cluster together while pushing those from different languages further apart in the embedding space. Combined with a classic CE loss, our dual-objective framework promotes learning less biased language representations that are robust to in-language domain shifts, enabling better generalization across diverse text types. A critical component of our SCL training method is a memory bank Tan et al. (2022) from which we scalably sample more diverse positive and negative examples, effectively increasing our training batch size while keeping memory requirements constant. This augmentation enhances our contrastive learning approach, yielding stable training and discriminative language representations. To mitigate the second challenge, we employ a novel hard negative mining scheme for SCL training, where data sources in our training set are labeled with domain information and negative samples are drawn from different languages in the same domain to learn language-specific, domain-invariant representations during training.
@@
We evaluate our approach on three benchmark datasets: GlotLID-C (Kargaran et al., 2023), FLORES-200 Team et al. (2022), and UDHR, [^2] https://huggingface.co/datasets/cis-lmu/udhr-lid and demonstrate that our approach leads to improvements over state-of-the-art LID methods, particularly in out-of-domain evaluations. Specifically, our best SCL-based approaches achieve a 3.2 percentage points improvement over CE-based models for low-resource languages, and a 5.4 percentage points improvement for languages with data from diverse domains, indicating enhanced domain generalization.
@@
Moreover, we conduct an in-depth analysis of underperforming languages in out-of-domain evaluations and observe that misclassifications predominantly occur among linguistically related languages, highlighting potential challenges in LID for closely related language pairs. Finally, we assess our approach in a real-world scenario by evaluating it on FineWeb-2 (Penedo et al., 2025), a large-scale multilingual pretraining corpus. We compare its performance against state-of-the-art LID systems for low-resource languages, providing insights into the practical significance of LID systems.
@@ 3 Methodology
In the following, we provide a detailed description of our base model architecture (§3.1), the principles of supervised contrastive learning (SCL; §3.2), our positive/negative pair selection approach (§3.3), and an optimized training strategy for the SCL objective, particularly suited for scenarios involving a large number of classes (§3.4).
@@ 3.1 Base Classifier
We train the base model following the methodology employed by FastText (Joulin et al., 2017). Given an input sentence, each word in the sentence is decomposed into character-level n-grams. The sentence representation is, then, obtained by computing the average of the vectors corresponding to these n-grams, along with the word-level embedding. The training process proceeds as follows:
Initially, the training dataset is partitioned into mini-batches. For a sampled mini-batch of size $B$, we define $S=\left\{x_{i},y_{i}\right\}_{i=1}^{B}$, where $x_{i}$ represents the input text, and $y_{i}$ denotes its corresponding language class. Each input text $x_{i}$ is represented by an embedding, $\mathbf{z}_{i}$, using an encoder function $\mathbf{z}_{i}=E(x_{i})$, where $E$ corresponds to the encoder. The resulting embeddings $\mathbf{z}_{i}$ are subsequently passed through a feed-forward neural network $f$, which serves as the classification head. The model is trained using the cross-entropy (CE) loss, formulated as:
@@ 3.3 Negative Pairs Selection
Given a batch $S=\left\{x_{i},y_{i};s_{i},d_{i}\right\}_{i=1}^{B}$, where $s_{i}$ and $d_{i}$ denote the script and domain of $x_{i}$, negative pairs are chosen as follows: for each sample $x_{i}$ in the batch, we select negative pair samples with a different label $y$, the same script $s$, and the same domain $d$. If fewer than $K$ such negatives exist, we relax the conditions in the following order: (2) different label $y$ and same script $s$, (3) different label $y$ and same domain $d$, (4) different label $y$ only (i.e., equivalent to Soft Selection). The algorithm of hard selection method is given in Appendix C.
@@ 4.1 Training
(a) $\text{LID}_{\text{CE}}$ - The model trained only with the cross-entropy objective, as defined in Equation (1).
(b) $\text{LID}_{\text{SCL}}$ - The model trained with the combined objective in Equation (3), without employing a memory bank.
@@
In our experiments, we implemented $\text{LID}_{\text{CE}}$ as a replication of GlotLID-M Kargaran et al. (2023), as official data splits for GlotLID-M were not openly released (hindering a non-contaminated evaluation on GlotLID-C-test). However, to enable a full comparison with prior methods, we also report the performance of the open-weight GlotLID-M, the previous SOTA model on UDHR, and GlotLID-M+ConLID-S, our ensemble of GlotLID-M and ConLID-S by taking the sum of their predicted distributions. Technical details can be found in Appendix B.1.
We also compare ConLID-S against two other publicly available LID models: AfroLID (Adebara et al., 2022) and NLLB-LID (Team et al., 2022).
@@ 4.2 Evaluation
$\text{LID}_{\text{CE}}$ .98586e-4 .00000681.2e-7 .892519e-4 .000222234e-7 .96503e-4 .00011429.9e-7
$\text{LID}_{\text{SCL}}$ .98331e-4 .00000750.2e-7 .893819e-4 .000209518e-7 .96173e-4 .000118225e-7
@@
The original UDHR dataset contains over 500 translations of the Universal Declaration of Human Rights. Kargaran et al. (2023) curated and refined this dataset, reducing it to 369 languages. During our experiments, we identified discrepancies in seven languages (dip_Latn, pcd_Latn, aii_Syrc, guu_Latn, mto_Latn, chj_Latn, auc_Latn) due to character misalignments between GlotLID-C and UDHR, and two languages (taj_Deva, tzm_Tfng) that had only one example each. We exclude these nine languages, resulting in a final UDHR dataset comprising 360 languages. UDHR serves as an out-of-domain evaluation dataset and is not included in GlotLID-C-train-down.
@@
This dataset is a multilingual pretraining corpus encompassing $\sim$1,800 languages (Penedo et al., 2025). The LID process in the FineWeb-2 pipeline is conducted using GlotLID-M Kargaran et al. (2023), a model capable of recognizing 2102 languages. We use this dataset to evaluate our model in large-scale real-world scenarios, which provides insights into the practical significance of our approach for LID for web crawls.
@@
[…]
High-resource 330 0.9065 0 0.0066 0.0066
Low-resource 30 0.7380 0.0168 0.0323 0.0154
[…]
Random 7 0.8157 0.0109 0.0541 0.0329
Bible 95 0.7821 -0.0077 0.0190 0.0149
[…]
@@ 5 Experimental Analysis
Table 1 presents the overall results for GlotLID-C-test, UDHR, and FLORES-200. We observe no significant improvement when directly applying the SCL objective ($\text{LID}_{\text{SCL}}$) compared to the cross-entropy baseline ($\text{LID}_{\text{CE}}$). However, when incorporating a memory bank to increase the number of contrastive negatives beyond the number of classes, we achieve performance gains across all three evaluation datasets. Notably, this improvement is most significant for UDHR (out-of-domain), indicating a better generalization of the contrastive learning approach.
For in-domain evaluation (GlotLID-C-test and FLORES-200), the highest F1 score among our methods is an ensemble approach that combines $\text{LID}_{\text{CE}}$ and ConLID-S with a memory bank of $\text{M}_{\text{H}}$=2048 and K=1024, suggesting that LIDCE already performs well for certain languages (as we will show in §5.1), while ConLID-S+LIDCE balances the performance across all labels.
When comparing to prior work, we observe that GlotLID-M tends to outperform our methods on FLORES-200, but underperforms them on UDHR, which is slightly more out of distribution. The ensembled method (GlotLID-M+ConLID-S), however, achieves the best score across UDHR and FLORES-200, indicating the methods are complementary and could be combined for optimal accuracy. No comparison is made on GlotLID-C-test as GlotLID-M was trained on different subsets of GlotLID-C, meaning the evaluation would be contaminated (same for GlotLID-M+ConLID-S).
Table 4 presents a performance comparison with AfroLID and NLLB-LID, evaluated on the subset of languages covered by all models. Results indicate that ConLID-S consistently surpasses both models across all three benchmark datasets.
Method GlotLID-C test UDHR FLORES-200
F1↑ FPR↓ F1↑ FPR↓ F1↑ FPR↓
ConLID-Sa .9860 .0000058 .9012 .0002170 .9652 .0001177
AfroLIDa .8609 .0002883 .7605 .0012766 .7618 .0011196
ConLID-Sb .9841 .0000058 .9639 .0002170 .9861 .0001177
NLLB-LIDb .9335 .0002380 .9405 .0003649 .9614 .0002708
Table 4: Comparison of ConLID-S with AfroLID and NLLB-LID systems. aEvaluated on 398/82/44 languages for GlotLID-C test/UDHR/FLORES-200 respectively. bEvaluated on 207/162/182 languages for GlotLID-C test/UDHR/FLORES-200 respectively.
@@ 5.1 Performance on UDHR Dataset
Resource Analysis.
As the performance of a language model is strongly influenced by the amount of training data available, we categorize languages into high- and low-resource groups based on the number of training examples available in GlotLID-C-train-down. Languages with fewer than 10k examples are designated as low-resource, while those with 10k or more examples are categorized as high-resource.
For each group of languages, we compute the average F1 score difference between SCL-based models ($\text{LID}_{\text{SCL}}$, ConLID-S, ConLID-S+LIDCE) and $\text{LID}_{\text{CE}}$. The results are presented in Table 2. Our analysis indicates that the model ConLID-S yields the most notable improvements for low-resource languages, with F1 score gain of 3.23 percentage points, validating the importance of memory bank (§3.4). In contrast, the performance gain for high-resource languages is more modest, at 0.66 percentage points. We further extend this analysis in Appendix D.3, where we present detailed results based on the low- and high-resource categorization defined by Joshi et al. (2020b).
Domain Analysis.
The GlotLID-C dataset comprises five distinct textual domains: Bible Translations, Literature (i.e., stories & books), Politics (i.e., government documents), Multi-domain (News and Wikipedia articles), and Random (machine translation datasets, chat logs, and web-scraped data). To evaluate the impact of SCL across different training data domains, we analyze languages trained exclusively on a single domain. While the distribution of GlotLID-C-train-down is shown in Appendix A, among the languages in UDHR, 7 languages have been exclusively trained on Random, while 95 on Bible. Therefore, we compute the F1 score difference between SCL-based models ($\text{LID}_{\text{SCL}}$, ConLID-S, ConLID-S+LIDCE) and $\text{LID}_{\text{CE}}$ for random and bible domains and present the results in Table 3. Our results indicate that the inclusion of a memory bank leads to performance improvements across both domains. Notably, the Random datasets exhibit a more significant gain, with improvements of up to 5.41 percentage points. We hypothesize that this is due to the broader range of subdomains covered by Random, which facilitate learning more generalized language representations. In contrast, the Bible domain is limited to religious texts, making it less conducive to generalization.
Script Analysis.
[…]
@@
[…]
5.2 Out-of-Domain Generalization
As noted GlotLID-C-train-down contains the following 5 domains: Bible, Literature, Politics, Multi-domain, Random. To further evaluate the out-of-domain generalization capabilities of our approach, we conducted the following experiment: we retained only the Bible domain from GlotLID-C-train-down and excluded the others. Using this single-domain dataset, we trained two models — one with SCL (ConLID-S-B) and one with CE ($\text{LID}_{\text{CE}}-\text{B}$). We then evaluated both models on all five domains of GlotLID-C, as well as on Flores-200 and UDHR. The resulting macro F1-scores are presented in Table 18.
The results demonstrate that even when training is limited to a single domain (i.e., Bible), our method improves generalization to previously unseen domains. Although the gains are modest under single-domain training, the advantage becomes substantially more pronounced when the model is trained on multiple domains.